# The Missing Point in Vision Transformers for Universal Image Segmentation

### v1

@@ 4.1 Datasets and Evaluation Metrics
Datasets. We evaluated our model using widely recognized image segmentation datasets that cover semantic, instance, and panoptic segmentation tasks. These datasets include COCO [31], which comprises 80 "things" and 53 "stuff" categories, ADE20K [56] with 100 "things" and 50 "stuff" categories, and Cityscapes [16] featuring 8 "things" and 11 "stuff" categories. Additionally, COCO-Stuff-164K [3], with 171 classes, was used for semantic segmentation. Panoptic and semantic segmentation tasks were assessed on the combined "things" and "stuff" categories, while instance segmentation focused exclusively on the "things" categories. Our comprehensive evaluation highlights the robustness of our model across diverse segmentation challenges.
Evaluation Metrics. To evaluate our model, we employ standard metrics across all three segmentation tasks. For panoptic segmentation, we use the PQ (Panoptic Quality) metric [29], which evaluates both recognition and segmentation quality. For instance segmentation, the standard AP (Average Precision) metric [31] is used, while semantic segmentation performance is measured using mIoU (mean Intersection-over-Union) [19].
@@ 4.2 Implementation Details
ViT-P is designed to be compatible with any pre-trained vision transformer backbone. In this study, we choose DINOv2 [36] for its strong performance and high-quality results. As a mask proposal generator, we employ OneFormer [27] as a universal model, while InternImage [46] serves as a non-universal alternative. During training, the mask proposal generator model remains frozen. For large backbone mask generation models, we use DINOv2 Base [36], while for the huge mask generator models, DINOv2 Large [36] is used. The number of input points is set to match the number of mask proposals.
@@
Random points within the input image are selected for classification during training. For inference, the points with the highest confidence scores in the generated masks are used as input to the ViT-P model for mask classification. We train the model using SGD optimizer with a warmup phase and a cosine learning rate scheduler. The training process consists of 30 epochs on COCO [31] and 60 epochs on ADE20K [56] and Cityscapes [16], with all setups incorporating 1000 warmup steps followed by a cosine scheduler and a learning rate of . Additionally, we found it beneficial to apply gradient clipping with a global norm of 1.
@@
We adopt a crop size of for COCO [31] and ADE20K [56] datasets and for Cityscapes [16]. Training is conducted on four A100 (40GB) GPUs, each utilizing a batch size of 20. Mixed precision floating point is used to increase efficiency. In the inference, we set .
Method #Params #FLOPs PQ AP UPerNet [48] SwinV2-L [32] — — 640640 — — — 55.9 UPerNet [48] InternImage-L [46] 256 2526 640640 — — 53.9 54.1 UPerNet [48] InternImage-XL [46] 368 3142 640640 — — 55.0 55.3 MaskFormer [15] Swin-L [33] 212 375 640640 — — 54.1 55.6 Mask2Former-Panoptic [14] Swin-L [33] 216 413 640640 48.7 34.2 54.5 — OneFormer [27] ConvNeXt-L [34] 220 389 640640 50.0 36.2 56.6 57.4 OneFormer [27] DiNAT-L [23] 223 1768 12801280 51.5 37.1 58.3 58.7 OneFormer† [27] DiNAT-L [23] 223 1768 12801280 53.4 40.2 58.4 58.8 OneFormer [27] ConvNeXt-XL [34] 372 607 640640 50.1 36.3 57.4 58.8 OpenSeeD†‡ [55] Swin-L [33] 286 — 12801280 53.7 42.6 58.4 — OneFormer+ViT-P (ours) DiNAT-L [23] 309 1955 12801280 51.9 37.8 58.6 59.0 OneFormer†+ViT-P (ours) DiNAT-L [23] 309 1955 12801280 54.0 40.7 59.7 59.9 UPerNet [48] SwinV2-G [32] 3 — 640640 — — 59.1 — OneFormer [27] InternImage-H [46] 1.1 4193 896896 54.5 40.2 60.4 60.8 Mask2Former† [14] RevCol-H [4] 2.4 — 640640 — — 60.4 61.0 Mask2Former† [14]+ViT-Adapter [11] BEiT-3 [45] 1.9 — 896896 — — 62.0 62.8 Mask2Former† [14] InternImage-H [46] 1.31 4635 896896 — — 62.5 62.9 OneFormer+ViT-P (ours) InternImage-H [46] 1.4 4812 896896 54.5 40.6 61.2 61.6 Mask2Former†+ViT-P† (ours) InternImage-H [46] 1.61 5254 896896 — — 63.1 63.6
Method Coarse #Params #Queries PQ AP Axial-DeepLab-XL [44] Axial ResNet-XL [44] 173 — 10252049 64.4 36.7 80.6 81.1 Panoptic-DeepLab [13] SWideRNet [10] 536 — 10252049 66.4 40.1 82.2 82.9 Mask2Former-Panoptic [14] Swin-L [33] 216 200 5121024 66.6 43.6 82.9 — Mask2Former-Semantic [14] Swin-L [33] 215 100 5121024 — — 83.3 84.3 OneFormer [27] ConvNeXt-L [34] 220 250 5121024 68.5 46.5 83.0 84.0 OneFormer† [27] ConvNeXt-L [34] 220 250 5121024 70.0 48.7 84.6 85.2 OneFormer† [27] ConvNeXt-XL [34] 372 250 5121024 69.7 48.9 84.5 85.8 OpenSeeD‡ [55] Swin-L [33] 286 300 5121024 68.9 48.5 84.5 — OneFormer†+ViT-P ConvNeXt-L [34] ✓ 306 250 5121024 70.0 49.0 84.9 85.5 OneFormer [27] InternImage-H [46] 1.1 250 5121024 70.6 50.6 85.1 85.7 Mask2Former [14]+Depth Anything [52] ViT-L [18] — 200 896896 — — 84.8 86.2 Mask2Former† [14] InternImage-H [46] 1.1 100 10241024 — — 86.4 87.0 OneFormer+ViT-P InternImage-H [46] ✓ 1.4 250 5121024 70.8 50.6 85.4 85.9 Mask2Former†+ViT-P† InternImage-H [46] ✓ 1.4 100 10241024 — — 86.8 87.4
COCO val2017 COCO-Stuff-164K Method #Params PQ AP MaskFormer [15] Swin-L [33] 212 640640 52.7 — 64.8 — — Mask2Former-Panoptic [14] Swin-L [33] 216 10241024 57.8 48.6 67.4 — — kMaX-DeepLab [54] ConvNeXt-L [34] 232 12811281 58.0 — — — — OneFormer [27] Swin-L [33] 219 10241024 57.9 49.0 67.4 — — OneFormer [27] DiNAT-L [23] 223 10241024 58.0 49.2 68.1 — — OpenSeeD† [55] Swin-L [33] 286 10241024 59.5 53.2 68.6 — — OneFormer+ViT-P (ours) DiNAT-L [23] 309 10241024 58.0 49.5 68.6 — — OneFormer+ViT-P† (ours) DiNAT-L [23] 309 10241024 58.2 49.6 68.8 — — OneFormer [27] InternImage-H [46] 1.35 10241024 60.0 52.0 68.8 — — Mask2Former [14]+ViT-Adapter [11] BEiTv2 [37] 571 896896 — — — 52.3 — Mask2Former [14]+ViT-Adapter [11] EVA [20] — 896896 — — — 53.4 — Mask2Former [14] InternImage-H [46] 1.31 896896 — — — 52.6 52.8 Mask2Former+ViT-P (ours) InternImage-H [46] 1.61 896896 — — — 53.5 53.7
4.3 Main Results
ADE20K. Without additional pre-training, the ViT-P model exhibits strong adaptability for mask classification on the ADE20K [56] dataset. As shown in Table 1, our approach is effective across all segmentation tasks, with the greatest impact on semantic segmentation. Using OneFormer [27] as the mask proposal method, our model achieves improvements, including a +0.6% increase in PQ, +0.5% in AP, and +1.3% in mIoU. Our proposed method reaches 63.6% mIoU on ADE20K[56], establishing a new state-of-the-art benchmark.
Cityscapes. The Cityscapes [16] dataset serves as a great benchmark for evaluating the effects of combining fine and coarse annotations, as it provides both annotation types. We conducted experiments assessing ViT-P’s performance under different settings and comparing it with the existing state-of-the-art architectures. Table 2 presents results on the Cityscapes [16] validation set, achieving an improvement of +0.2% in PQ, +0.3% in AP, and +0.5% in mIoU metrics. With InternImage [46] as the mask generation backbone, our model sets a new benchmark, achieving 87.4% mIoU.
COCO. We evaluate ViT-P against state-of-the-art universal and specialized architectures on the COCO [31] val2017 and COCO-Stuff-164K [3] datasets in Table 3. ViT-P achieves 68.6% mIoU on COCO [31] and 53.5% on COCO-Stuff-164K [3]. Additionally, inspired by Mask-DINO [30] and OpenSeeD [55], which leverage box annotations for improved segmentation, we pre-train ViT-P using the Objects365 [40] dataset. In this approach, random points are selected within each bounding box and assigned the corresponding label. This method enables ViT-P to achieve state-of-the-art performance, reaching 56.8% mIoU on COCO [31], with slight improvements of +0.2% for instance segmentation and +0.4% for panoptic segmentation in OneFormer [27].
4.4 Ablation Studies
We conducted an in-depth analysis of the ViT-P structure through various ablation studies. Unless otherwise specified, all ablations were performed with the OneFormer [27] mask generator on the ADE20K [56] dataset.
Number of Input Points. Adjusting the number of input points in the ViT-P model affects both mask proposal classification and overall training performance. Table 7 shows the results for different input configurations. Insufficient points lead to poor training quality. However, beyond a certain threshold, adding more points does not enhance classification accuracy.
@@ 4.4 Ablation Studies
Inference Mask Classification Point. By analyzing the points inside each mask, the ViT-P model classifies mask proposals. In Table 7, we evaluate three methods for selecting points within a mask: random selection, the central point (the farthest from mask boundaries), and the highest-value point. Random selection results in a lower mIoU of 58.7%, while both the central and highest-value points improve performance by 1.0% mIoU. The highest-value point is preferred as it requires no additional computation and is typically near the mask center.
Classification Backbone. To evaluate the adaptability of the ViT-P model, we conducte an ablation study with four pre-trained vision transformer backbones: a plain ViT [18], DinoV2 [36], DinoV2 with additional registers (DinoV2+reg) [17], and MAE [24] as shown in Table 7. While the plain ViT [18] remains effective and enhances mask classification performance, the DinoV2 [36] backbone exhibits better compatibility, yielding superior results. Notably, since the vision transformer architecture in the ViT-P model remains unchanged, it can be replaced with more advanced pre-trained transformers in the future, further enhancing its mask classification performance.
@@ 5 Conclusion
In this paper, we proposed ViT-P, a universal image mask classification model based on vision transformers, addressing the limitations of traditional mask-based segmentation methods. By introducing a two-stage approach, ViT-P separates mask generation from classification, using a proposal generator and a point-based classification model. This method enhances the accuracy of classification while maintaining a high-quality mask generation. Our comprehensive experiments demonstrate that ViT-P outperforms existing state-of-the-art segmentation models across multiple benchmarks, achieving new state-of-the-art results on COCO, ADE20K, and Cityscapes datasets. The main advantage of ViT-P is its ability to leverage pre-trained weights while maintaining minimal additional layers, making it both efficient and adaptable. This design bridges the gap between standard vision transformers and specialized models for dense prediction tasks. This approach can be further refined to enhance mask classification accuracy, progressively working toward the upper bound, while also improving mask quality to ensure more precise and reliable segmentation.
Appendix A Appendix
A.1 Upper Bound Analysis
@@
A.2 Mask-Prompting vs. Point-Prompting
While the ViT-P model classifies mask proposals based on input points, an alternative approach is to directly use mask proposals as inputs into a classification model. Prior works, such as MAFT [28], explored this approach by feeding the entire mask into a transformer model for labeling. In this section, we compare mask-based prompting with point-based prompting and explain why we favor point-based classification over using the complete mask proposal. To ensure a fair evaluation, both the ViT-P and MAFT [28] models were trained on the ADE20K [56] dataset using two different backbones, CLIP [39] and DinoV2 [36], with the Mask2Former [14] model serving as the mask generator.
As shown in Table 9, point-based classification, ViT-P, outperforms mask-based prompting, MAFT [28], across all three segmentation tasks. By adding only a linear layer to the pre-trained transformer backbone, ViT-P preserves the original architecture. In contrast, the MAFT [28] approach integrates additional randomly initialized attention layers into the -layer backbone. This extra layer stack not only increases complexity but also limits the model’s adaptability when paired with standard vision transformer models such as CLIP [39] and DinoV2 [36] for dense prediction tasks, ultimately decreasing its performance.
In addition, mask-prompting models such as MAFT [28] suffer from mask quality during training. Typically, the number of generated masks exceeds the available ground truth masks, leading to the inclusion of additional low-quality mask proposals. Training on these generated masks reduces model accuracy, as overlapping masks introduce noise that negatively impacts learning efficiency. Conversely, training on high-quality ground truth masks creates an inconsistency between the training and inference, as the model is ultimately evaluated on imperfectly generated mask proposals during testing. These inaccuracies introduce inconsistencies that degrade classification performance. In contrast, the ViT-P model benefits from stable input points throughout both training and evaluation, ensuring more consistency. Theses findings highlight the limitations of mask-based prompting relative to point-based classification. Consequently, providing a strong justification for our decision to adopt a point-based approach in this study.
A.3 Limitations
@@
A.4 Qualitative Results
We present segmentation examples on three benchmark datasets, ADE20K [56], COCO [31], and Cityscapes [16], illustrating ViT-P model’s performance, in Figure 4, 5, 6.

### v2

@@ 4.1 Datasets and Evaluation Metrics
Datasets. We evaluated our model using widely recognized image segmentation datasets that cover semantic, instance, and panoptic segmentation tasks. These datasets include COCO [lin2014microsoft], which comprises 80 ”things” and 53 ”stuff” categories, ADE20K [zhou2017scene] with 100 ”things” and 50 ”stuff” categories, and Cityscapes [cordts2016cityscapes] featuring 8 ”things” and 11 ”stuff” categories. Additionally, COCO-Stuff-164K [caesar2018coco], with 171 classes, was used for semantic segmentation. Panoptic and semantic segmentation tasks were assessed on the combined ”things” and ”stuff” categories, while instance segmentation focused exclusively on the ”things” categories. Our comprehensive evaluation highlights the robustness of our model across diverse segmentation challenges.
Evaluation Metrics. To evaluate our model, we employ standard metrics across all three segmentation tasks. For panoptic segmentation, we use the PQ (Panoptic Quality) metric [kirillov2019panoptic], which evaluates both recognition and segmentation quality. For instance segmentation, the standard AP (Average Precision) metric [lin2014microsoft] is used, while semantic segmentation performance is measured using mIoU (mean Intersection-over-Union) [everingham2015pascal].
@@ 4.2 Implementation Details
ViT-P is designed to be compatible with any pre-trained vision transformer backbone. In this study, we choose DINOv2 [oquab2023dinov2] for its strong performance and high-quality results. As a mask proposal generator, we employ OneFormer [jain2023oneformer] as a universal model, while InternImage [wang2023internimage] serves as a non-universal alternative. During training, the mask proposal generator model remains frozen. For large backbone mask generation models, we use DINOv2 Base [oquab2023dinov2], while for the huge mask generator models, DINOv2 Large [oquab2023dinov2] is used. The number of input points is set to match the number of mask proposals.
@@
Random points within the input image are selected for classification during training. For inference, the points with the highest confidence scores in the generated masks are used as input to the ViT-P model for mask classification. We train the model using SGD optimizer with a warmup phase and a cosine learning rate scheduler. The training process consists of 30 epochs on COCO [lin2014microsoft] and 60 epochs on ADE20K [zhou2017scene] and Cityscapes [cordts2016cityscapes], with all setups incorporating 1000 warmup steps followed by a cosine scheduler and a learning rate of . Additionally, we found it beneficial to apply gradient clipping with a global norm of 1.
@@
We adopt a crop size of for COCO [lin2014microsoft] and ADE20K [zhou2017scene] datasets and for Cityscapes [cordts2016cityscapes]. Training is conducted on four A100 (40GB) GPUs, each utilizing a batch size of 20. Mixed precision floating point is used to increase efficiency. In the inference, we set .
@@
PQ AP mIoU DINOv2 [oquab2023dinov2] 54.0 40.7 59.7 ViT [dosovitskiy2020image] 53.4 40.2 58.8 DINOv2+reg [darcet2023vision] 53.7 40.6 59.3 MAE [he2022masked] 53.6 40.6 59.1
@@
4.3 Main Results
ADE20K. Without additional pre-training, the ViT-P model exhibits strong adaptability for mask classification on the ADE20K [zhou2017scene] dataset. As shown in Table 1, our approach is effective across all segmentation tasks, with the greatest impact on semantic segmentation. Using OneFormer [jain2023oneformer] as the mask proposal method, our model achieves improvements, including a +0.6% increase in PQ, +0.5% in AP, and +1.3% in mIoU. Our proposed method reaches 63.6% mIoU on ADE20K[zhou2017scene], establishing a new state-of-the-art benchmark.
Cityscapes. The Cityscapes [cordts2016cityscapes] dataset serves as a great benchmark for evaluating the effects of combining fine and coarse annotations, as it provides both annotation types. We conducted experiments assessing ViT-P’s performance under different settings and comparing it with the existing state-of-the-art architectures. Table 2 presents results on the Cityscapes [cordts2016cityscapes] validation set, achieving an improvement of +0.2% in PQ, +0.3% in AP, and +0.5% in mIoU metrics. With InternImage [wang2023internimage] as the mask generation backbone, our model sets a new benchmark, achieving 87.4% mIoU.
COCO. We evaluate ViT-P against state-of-the-art universal and specialized architectures on the COCO [lin2014microsoft] val2017 and COCO-Stuff-164K [caesar2018coco] datasets in Table 3. ViT-P achieves 68.6% mIoU on COCO [lin2014microsoft] and 53.5% on COCO-Stuff-164K [caesar2018coco]. Additionally, inspired by Mask-DINO [li2023mask] and OpenSeeD [zhang2023simple], which leverage box annotations for improved segmentation, we pre-train ViT-P using the Objects365 [shao2019objects365] dataset. This method enables ViT-P to achieve state-of-the-art performance, reaching 56.8% mIoU on COCO [lin2014microsoft], with slight improvements of +0.2% for instance segmentation and +0.4% for panoptic segmentation in OneFormer [jain2023oneformer].
4.4 Ablation Studies
We conducted an in-depth analysis of the ViT-P structure through various ablation studies. Unless otherwise specified, all ablations were performed with the OneFormer [jain2023oneformer] mask generator on the ADE20K [zhou2017scene] dataset.
Number of Input Points. Adjusting the number of input points in the ViT-P model affects both mask proposal classification and overall training performance. Table 4(a) shows the results for different input configurations. Insufficient points lead to poor training quality. However, beyond a certain threshold, adding more points does not enhance classification accuracy.
Inference Mask Classification Point. By analyzing the points inside each mask, the ViT-P model classifies mask proposals. In Table 4(b), we evaluate three methods for selecting points within a mask: random selection, the central point (the farthest from mask boundaries), and the highest-value point. Random selection results in a lower mIoU of 58.7%, while both the central and highest-value points improve performance by 1.0% mIoU. The highest-value point is preferred as it requires no additional computation and is typically near the mask center.
Classification Backbone. To evaluate the adaptability of the ViT-P model, we conducte an ablation study with four pre-trained vision transformer backbones: a plain ViT [dosovitskiy2020image], DinoV2 [oquab2023dinov2], DinoV2 with additional registers (DinoV2+reg) [darcet2023vision], and MAE [he2022masked] as shown in Table 4(c). While the plain ViT [dosovitskiy2020image] remains effective and enhances mask classification performance, the DinoV2 [oquab2023dinov2] backbone exhibits better compatibility, yielding superior results. Notably, since the vision transformer architecture in the ViT-P model remains unchanged, it can be replaced with more advanced pre-trained transformers in the future, further enhancing its mask classification performance.
Coarse Annotations Impact. Training the ViT-P model on Cityscapes [cordts2016cityscapes] with combined fine and coarse annotations results in additional accuracy improvements, Table 4(d). Integrating these annotations offers an effective strategy for better segmentation performance.
@@ 5 Conclusion
In this paper, we proposed ViT-P, a universal image mask classification model based on vision transformers, addressing the limitations of traditional mask-based segmentation methods. By introducing a two-stage approach, ViT-P separates mask generation from classification, using a proposal generator and a point-based classification model. This method enhances the accuracy of classification while maintaining a high-quality mask generation. Our comprehensive experiments demonstrate that ViT-P outperforms existing state-of-the-art segmentation models across multiple benchmarks, achieving new state-of-the-art results on COCO, ADE20K, and Cityscapes datasets. The main advantage of ViT-P is its ability to leverage pre-trained weights while maintaining minimal additional layers, making it efficient and adaptable. This design bridges the gap between standard vision transformers and specialized models for dense prediction tasks. This approach can be further refined to enhance mask classification accuracy, progressively working toward the upper bound, while also improving mask quality to ensure more precise and reliable segmentation.
6 Upper Bound Analysis
@@
Model Backbone Classification Method OneFormer [jain2023oneformer] DiNAT-L [hassani2022dilated] Point Labeling 86.5 88.7 83.5 Mask Labeling 87.3 89.7 84.7
7 Mask-Prompting vs. Point-Prompting
While the ViT-P model classifies mask proposals based on input points, an alternative approach is to directly use mask proposals as inputs into a classification model. Prior works, such as MAFT [jiao2023learning], explored this approach by feeding the entire mask into a transformer model for labeling. In this section, we compare mask-based prompting with point-based prompting and explain why we favor point-based classification over using the complete mask proposal. To ensure a fair evaluation, both the ViT-P and MAFT [jiao2023learning] models were trained on the ADE20K [zhou2017scene] dataset using two different backbones, CLIP [radford2021learning] and DinoV2 [oquab2023dinov2], with the Mask2Former [cheng2022masked] model serving as the mask generator.
As shown in Table 6, point-based classification, ViT-P, outperforms mask-based prompting, MAFT [jiao2023learning], across all three segmentation tasks. By adding only a linear layer to the pre-trained transformer backbone, ViT-P preserves the original architecture. In contrast, the MAFT [jiao2023learning] approach integrates additional randomly initialized attention layers into the -layer backbone. This extra layer stack not only increases complexity but also limits the model’s adaptability when paired with standard vision transformer models such as CLIP [radford2021learning] and DinoV2 [oquab2023dinov2] for dense prediction tasks, ultimately decreasing its performance.
In addition, mask-prompting models such as MAFT [jiao2023learning] suffer from mask quality during training. Typically, the number of generated masks exceeds the available ground truth masks, leading to the inclusion of additional low-quality mask proposals. Training on these generated masks reduces model accuracy, as overlapping masks introduce noise that negatively impacts learning efficiency. Conversely, training on high-quality ground truth masks creates an inconsistency between the training and inference, as the model is ultimately evaluated on imperfectly generated mask proposals during testing. These inaccuracies introduce inconsistencies that degrade classification performance. In contrast, the ViT-P model benefits from stable input points throughout both training and evaluation, ensuring more consistency. Theses findings highlight the limitations of mask-based prompting relative to point-based classification. Consequently, providing a strong justification for our decision to adopt a point-based approach in this study.
Method Classification Backbone PQ AP mIoU Mask2Former [cheng2022masked] — 48.7 34.2 54.5 Mask2Former [cheng2022masked]+MAFT [jiao2023learning] CLIP [radford2021learning] 47.6 33.3 54.6 Mask2Former [cheng2022masked]+ViT-P 49.0 34.4 55.3 Mask2Former [cheng2022masked]+MAFT [jiao2023learning] DinoV2 [oquab2023dinov2] 48.6 34.2 54.8 Mask2Former [cheng2022masked]+ViT-P 49.3 34.7 56.1
8 Limitations
@@
9 Qualitative Results
We present segmentation examples on three benchmark datasets, ADE20K [zhou2017scene], COCO [lin2014microsoft], and Cityscapes [cordts2016cityscapes], illustrating ViT-P model’s performance, in Figure 4, 5, 6.